Tan75+tan15/1-tan75×tan15 please answer this question...I need the answer fast. Thanks whoever answers.

Grade:10

1 Answers

Shailendra Kumar Sharma
6 years ago
) We can relate it to (Tan75+tan15)/(1-tan75×tan15 (TanA+tanB)/(1-tanA×tanB) =tan(A+B)
So ) =Tan (75+15) =tan90 =infinity  (Tan75+tan15)/(1-tan75×tan15